### v1

# The cosmic trimmer: Black-hole hair in scalar-Gauss-Bonnet gravity is altered by cosmology

Abstract
Static black holes in general relativity modified by a linear scalar coupling to the Gauss-Bonnet invariant always carry hair. We show that the same mechanism that creates the hair makes it incompatible with a cosmological horizon. Other scalar-tensor models do not have this problem when time-dependence of the scalar provides a natural matching to cosmology. Scalar-Gauss-Bonnet is particularly rigid and such a scenario does not help. An extra operator makes the theory behave like the other models and the cosmological horizon can be accommodated. The hair, however, is drastically altered.
I Introduction
The detection of gravitational waves emitted from the final stages of the inspiral of binary systems of compact objects Abbott et al. (2016, 2017) has opened a new and finally direct avenue for testing the validity of general relativity (GR) as the theory of gravity. These tests, however, require alternative models of gravity which could be used as a laboratory for the prediction of realistic deviations from GR phenomenology. The existence of hair attached to scalar charges carried by compact objects in scalar-tensor theories of gravity provides just such a opportunity. This is the existence of alternative solutions for star-like objects in addition to or instead of the Schwarzschild or Kerr cases possible for the star’s exterior in GR. These hairy (“scalarised”) solutions were discovered in the case of conformally coupled theories of gravity, where it is only possible for neutron stars but not black holes (BH) Bocharova et al. (1970); Damour and Esposito-Farese (1993). Despite the no-hair theorems, see e.g. Bekenstein (1995); Hui and Nicolis (2013), even black holes can carry hair under certain circumstances, for example, when the theory of gravity includes interaction terms between the scalar and higher-order curvature corrections, such as the Gauss-Bonnet term (GB), in principle expected not in the least from effective field theory (EFT) arguments. This alternative phase usually exists only for ranges of BH parameters and appears as a (tachyonic) instability of the GR solution, and can lead to modifications of observable properties of the black holes, such as changes in the spectrum of quasi-normal oscillations or modifications of the BH exterior.
However, the embedding of such setups in cosmological backgrounds presents a challenge. Typically, the parameters which allow for hair in high-curvature environments while permitting a return to GR far from the objects, also cause the equivalent tachyonic instability to appear in the cosmological past. The result is a catastrophic modification of cosmological history compared to the one predicted from GR Babichev et al. (2022). This is incompatible with observations. Perhaps uniquely, the linear coupling to the GB term is not known to suffer from this instability. This theory is quite particular, in that the finiteness of physical observables on the BH horizon in fact requires that BHs carry a charge and the associated hair, making them different than in GR. This has led to a significant body of work studying this model, for a review see Babichev et al. (2016); Lehébel (2018); Herdeiro and Radu (2015) and references within.
Notwithstanding cosmological evolution, attempts have been made to find numerically a static BH solution in scalar-GB gravity in a de Sitter background. They have not been successful Bakopoulos et al. (2019) for so-far unclear reasons.
In this letter, we discuss the case of a linear scalar coupling to GB (sGB) in the context of a black hole embedded within a cosmology under the assumptions of static spacetime and negligible backreaction. We demonstrate that the same finiteness condition that creates the BH hair, makes the cosmological/de Sitter horizon incompatible with it, explaining why such solutions cannot be found. As a result of the shift-symmetry, we can introduce a linear time dependence for the scalar without affecting observables, which has the benefit of being a natural setup in the cosmological embedding. However, this does not solve the problem for this theory. A non-evolving solution in this setup cannot be found.
We show an explicit example of a shift-symmetric theory where this stationary setup does work. We propose that the problem in linear sGB can be solved by a small extension of the model with additional shift-symmetric operators together with the cosmologically induced time dependence. In this case, the horizons no longer provide any constraint on the value of the scalar charge. We expect that this new setup would introduce substantial changes to the dynamics of black holes when compared to the static solutions in linear sGB.
II Static hair in scalar-Gauss-Bonnet theory
It is instructive to first revisit the usual static hairy black-hole solutions in the standard linear sGB theory,
$S=M_{P}^{2}\int d^{4}\!x\sqrt{-g}\left[\frac{R}{2}-\frac{1}{2}(\partial\phi)^{2}+\alpha\phi\,\mathcal{G}\right]\,,$ (1)
where $\mathcal{G}=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2},$ is the Gauss-Bonnet invariant, $\phi$ is dimensionless, and $M_{P}$ is the Planck mass. The corresponding scalar field equation is
$\square\phi+\alpha\mathcal{G}$ $=$ $0\,,$ (2)
The size of the backreaction caused by the sGB term on the background geometry is dictated by the ratio $\alpha/L^{2}$, where $L$ is a characteristic curvature scale of the spacetime Sotiriou and Zhou (2014). Notice that, since sGB is a shift-symmetric Horndeski theory Kobayashi et al. (2011), the scalar field equation (2) can be written as
$\nabla_{\mu}J^{\mu}=0\,,$ (3)
where the contribution of the sGB term to the current $J^{\mu}$, while not unique in general Yale and Padmanabhan (2011) (see also Creminelli et al. (2020)), it can for example be obtained from the quintic Horndeski current with $G_{5}\sim\log(X)$.
For a static, spherically-symmetric background metric and scalar of the form
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2}\,,\qquad\phi=\phi(r)\,,$ (4)
the scalar field equation can easily be integrated to give
$J_{r}=-\phi^{\prime}+\frac{4\alpha f^{\prime}}{r^{2}f}(1-f)=\frac{C}{r^{2}f}\,,$ (5)
where $C$ is an integration constant to be determined. In order to fix the value of this constant, the standard procedure is to demand the finiteness of $\phi^{\prime}$ and $\phi^{\prime\prime}$ at the black-hole horizon, as proxies for observable quantities. Alternatively, a more covariant and unambiguous way to achieve the same is to require that the kinetic term $X=-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$ remain finite, as it appears for instance in the stress-energy tensor. Near a horizon this goes like
$X=-\frac{f\phi^{\prime 2}}{2}\simeq-\frac{1}{2fr_{h}^{4}}\left(C-4\alpha f^{\prime}_{h}\right)^{2}\,,$ (6)
and it remains finite only if
$C=4\alpha f^{\prime}_{h}\,.$ (7)
Under a test-field approximation, we can take the background metric to be the Schwarzschild one, $f=1-r_{s}/r$, with $r_{s}$ the Schwarzschild radius, and the backreaction being under control as long as $\alpha\ll r_{s}^{2}$. Then it is possible to integrate the scalar field to obtain Sotiriou and Zhou (2014)
$\phi=\frac{4\alpha}{r_{s}r}\left(1+\frac{r_{s}}{2r}+\frac{r_{s}^{2}}{3r^{2}}\right)\,,$ (8)
which asymptotically goes like $C/r$. The black holes in this theory therefore carry a secondary scalar charge, that is, it is not a free parameter but is instead fixed in terms of the black-hole mass $M=M_{P}^{2}r_{s}/2$, as given by Eq. (7). Importantly, it is the presence of a horizon that is responsible for the charge, while horizonless sources such as neutron stars can only support a much-faster-decaying scalar profile Yagi et al. (2016).
Beyond the test-field approximation above, perturbative corrections to the metric have been studied alongside fully self-consistent numerical solutions in Ref. Sotiriou and Zhou (2014), and also considering slow rotation Ayzenberg and Yunes (2014). The observed absence of dephasing on the gravitational-wave waveform during the inspiral phase of binary black-hole mergers has established the most stringent bound on the sGB coupling so far to be $\alpha\lesssim(1.2\,\text{km})^{2}$ Lyu et al. (2022). This justifies the perturbative treatment for $\alpha\ll r_{s}^{2}$ for astrophysical black holes. Further studies have also constrained the validity of the EFT (1), which must have a UV cutoff $\Lambda_{UV}<1/\sqrt{\alpha}$ in order to avoid resolvable superluminal propagation Serra et al. (2022).
Now let us extend the model to include a positive cosmological constant $\Lambda>0$, producing a still static but expanding spacetime with Hubble constant $H$. For observationally allowed values of $\alpha$, $H$ is completely driven by $\Lambda$ and the correction to the speed of gravitational waves is completely negligible. In the test-field approximation, we would expect the Schwarzschild-de Sitter GR solution for the spacetime to carry across, only modifying the metric function to
$f=1-\frac{r_{s}}{r}-H^{2}r^{2}\,.$ (9)
In the astrophysically relevant scenario $Hr_{s}\ll 1$, the Schwarzschild horizon remains at $r_{h}\simeq r_{s}$, but $f$ also has second root and therefore a second horizon – the standard Hubble horizon of cosmology located at $r_{h}\simeq H^{-1}$. For a scalar-field solution valid in the whole observable region, we also have to enforce the finiteness of $X$ on the new cosmological horizon. Eq. (7) then requires that $C$ simultaneously takes different values at each of the horizons, while being a constant and therefore is impossible to satisfy. This is the origin of the non-construction of black-hole solution in the presence of a positive $\Lambda$ in Ref. Bakopoulos et al. (2019). In the next section we will demonstrate that in cosmology the scalar field cannot remain static, and therefore the time dependence must be taken into account when attempting to solve the two-horizon problem.
III Time-dependent hair in cosmologically embedded black-holes
Cosmology suggests a generalisation of this setup which might resolve the problem: it is natural to consider a time-dependent scalar field. We take the simplest ansatz which contains both a time dependence and the spherically symmetric metric required by the local source,
$\phi=q\,t+\varphi(r)\,.$ (10)
$q$ is a new parameter which might allow both the horizons to be compatible with the solution. Owing to the invariance of the action (1) under shifts of $\phi$, such linear time dependence does not give rise to time dependence of any observable quantity. Thus it is compatible with static solutions in principle, as long as the effect of accretion can be neglected. As for the static case, it is again possible to integrate the scalar field equation Eq. (3) once, giving,
$J_{r}=-\varphi^{\prime}+\frac{4\alpha f^{\prime}}{r^{2}f}(1-f)=\frac{C}{r^{2}f}\,,$ (11)
where again $C$ is a constant of integration yet to be determined. Incidentally, this is the same equation as for the static case, up to the replacement $\phi\to\varphi$.
Following the same logic as for the static case, we will demand that the scalar quantity $X$ be finite at both horizons. With the ansatz above, the kinetic term now is
$X=\frac{1}{2}\left(\frac{q^{2}}{f}-f\varphi^{\prime 2}\right)\,,$ (12)
which implies that, unlike in the static case, $\varphi$ can no longer be finite itself if $X(r_{h})=X_{h}$ is to not diverge. Since no observable quantity depends on $\phi$, while $X$ is explicitly contained in at least the energy-momentum tensor, we demand only the finiteness of $X_{h}$ and parameterize the $r$-dependent part of the scalar near any horizon as
$\varphi^{\prime}$ $=$ $\epsilon\frac{q}{f}+\psi^{\prime}(r)\,,$ (13)
with $\epsilon(r_{h})=\pm 1$, and where $\psi(r_{h})$ is now manifestly finite and related to $X_{h}$ as
$X_{h}=-\epsilon_{h}q\,\psi^{\prime}(r_{h})\,.$ (14)
Then in this near-horizon limit the scalar field equation (11) multiplied by $f$ reads
$\frac{C}{r_{h}^{2}}=J^{r}|_{h}$ $=$ $-\epsilon_{h}q+4\alpha\frac{f^{\prime}_{h}}{r_{h}^{2}}+O(f)\,.$ (15)
Notice that at this order there is no dependence on $X_{h}$, or equivalently $\psi^{\prime}_{h}$. It can now easily be checked that the scalar quantity $J^{2}$ remains finite at the horizons for any value of $C$ [^1] This follows from the fact that, near the horizon, $J_{0}(r_{h})\simeq\pm J^{r}(r_{h})+O(f)$, and therefore $J^{2}\sim 1/q$, which is finite for $q\neq 0$. This is contrary to the static case where the divergence was unavoidable already at the BH horizon Babichev et al. (2016).
Now let us, in turn, examine each horizon and see which conditions arise on $C$. The condition on the black-hole horizon gives
$C$ $=$ $r_{s}^{2}\left(-\epsilon_{h}q+\frac{4\alpha}{r_{s}^{3}}\right)\,.$ (16)
Choosing $C$ requires that we pick some $q$ – this is the extra freedom we have introduced. We can do so by examining the cosmological horizon, Eq. (15) naively gives
$q$ $=$ $H^{2}C+8\alpha H^{3}+O(Hr_{s})\,,$ (17)
Seemingly we have found our new solution compatible with both horizons. The question now is how this connects to the cosmological solution.
Let us discuss the theory (1) in a cosmological setting in the absence of a local source. For simplicity we consider de Sitter space,
$ds^{2}=-d\tau^{2}+e^{2H\tau}d{\bf x}^{2}\,,$ (18)
where the scalar-field equation takes the simple form
$\ddot{\phi}+3H\dot{\phi}=24\alpha H^{4}\,,$ (19)
with ˙$\dot{}=\partial_{\tau}$. This equation admits an attractor ($\ddot{\phi}=0$) solution
$q_{c}\equiv\dot{\phi}=8\alpha H^{3}\,.$ (20)
Thus the cosmology imposes that the scalar field always asymptotes to a rolling stationary solution
$\phi_{c}=q_{c}\,\tau\,.$ (21)
Translating the cosmological coordinates to the static ones, this implies that in the $r\rightarrow\infty$ limit the cosmological $\phi$ tends to
$\phi=q_{c}\,t+\int\mathrm{d}r\varphi^{\prime}\,,\quad\varphi^{\prime}\simeq\frac{q_{c}}{Hr}\,.$ (22)
The only consistent possibility when embedding a BH in cosmology is that the solution sufficiently far away from the BH becomes homogeneous. We thus need to match the scalar hair to this cosmological boundary condition. Comparing this expression for $\varphi^{\prime}$ with the one implied by (11) for $r\rightarrow\infty$ we find that $q$ must take the cosmological value $q_{c}$, while the finiteness condition instead gives
$\frac{q}{q_{c}}=\frac{1}{2Hr_{s}}+1+O(Hr_{s})\gg 1\,.$ (23)
This is vastly different from the one required by cosmology. Moreover, this discrepancy is larger for small black holes. As a matter of principle, the local solutions should not affect the cosmology strongly, but rather the cosmology should represent the aggregated effect of the averaged localized sources – this principle is violated by the static solution in sGB.
This is the main result of this letter: There is an impossibility to construct black-hole solutions in the most simple sGB theory (1) with the proper cosmological asymptotics, when the metric is static and the scalar field time dependent at most in a manner which keeps observables static. The problem results from needing a large charge $C$ to keep the scalar quantities ($X$) finite at the black-hole horizon, which is then incompatible with the same requirement at the cosmological horizon.
One may think that the two-horizon problem could be fixed without much change to the expected physics by breaking the shift-symmetry and adding a potential $V(\phi)$ to the scalar field. In this case the scalar can be asymptotically static at the minimum of $V_{,\phi}+\alpha\mathcal{G}$. Picking $V(\phi)$ such that the mass is of inverse galactic scales would then screen the hair at cosmological distances while giving astrophysics as expected in the static shift-symmetric case. This is true – this modificaton would allow us to use the static solution in the completely static universe. However, during the approach to the asymptotic future de Sitter state, the minimum of the potential evolves and the scalar must roll slowly, giving a small $\dot{\phi}=q$. This would then be enough to put us in the scenario in which the finiteness of $X$ on the horizon would imply that $\phi$ diverges. Since $\phi$ is now a physical observable, this is inconsistent. It remains to be seen whether any such non-shift-symmetric resolution of the problem would not alter the nature of the standard static solution.
Next we investigate an alternative approach to solving the problem that maintains shift-symmetry.
IV The good example: Cubic Galileon
A fair question is whether the type of problem appears in other scalar-tensor theories. This turns out not to be the case in general. As an example of a model which is free of this particular issue we now consider the Cubic Galileon model discussed in Ref. Babichev and Esposito-Farèse (2013), with action
$S=M_{P}^{2}\int d^{4}\!x\sqrt{-g}\left[\frac{R}{2}+\frac{1}{2}(\partial\phi)^{2}+\frac{(\partial\phi)^{2}}{2\Lambda^{2}}\square\phi\right]\,.$ (24)
Notice that the kinetic term has a ghostly sign, which leads to self-accelerated cosmological solutions Deffayet et al. (2010). The homogeneous solution for the scalar of the form (21), asymptotes toward
$q_{c}=-\frac{\Lambda^{2}}{3H}\,.$ (25)
Now, turning on the mass of the black hole, we consider as previously the ansatz (10) for the Schwarzschild-de Sitter background metric. The analogue of (11) in this case is
$J_{r}=\varphi^{\prime}+\frac{1}{\Lambda^{2}}\left(\frac{Xf^{\prime}}{f}-\frac{2f\varphi^{\prime 2}}{r}\right)=\frac{C}{r^{2}f}\,.$ (26)
The above expression cannot be satisfied at all unless $X$ is finite, implying the solution of the form (13). This means that the value of $X$ is necessarily finite at a horizon for any $C$. This is in contrast to the case of sGB theory, where we had to fix $C$ to avoid the divergence of $X_{h}$. The value of $X_{h}$ at a horizon is found from (26) by taking the limit $r\to r_{h}$,
$\frac{C}{r_{h}^{2}}=\epsilon_{h}q-\frac{1}{\Lambda^{2}}\left(-X_{h}f_{h}^{\prime}+\frac{2q^{2}}{r_{h}}\right)+O(f)\,,$ (27)
Applying the same argument as below Eq. (22), also forces $q=q_{c}$ when homogeneity is required in the $r\to\infty$ limit. This is true for any $C$.
V Extra operators alleviate the problem
The fact that the example of the Cubic Galileon works well, in contrast to the sGB case, can be traced to the form of the current (26). The explicit presence of $X$ in this equation forces $X_{h}$ to be finite. In this respect, the sGB operator is rather unique among the Horndeski operators in not providing an explicit $X$ term in the current (11). The natural question is then whether the presence of any of these additional operators may mitigate the problem faced by the vanilla sGB theory (1).
To have an explicit example, it is not difficult to check that when combining the sGB term with the Cubic Galileon term, the solution for the scalar field remains of the form (13), thus providing a finite $X_{h}$ at both horizons. We expect similar results when we combine sGB with other Horndeski terms that contain a function of $X$. Indeed, assuming the behaviour (13), for the sGB theory with an extra shift-symmetric term, the equation of motion for the scalar at a horizon becomes (cf. Eq. (27)),
$\frac{C}{r_{h}^{2}}=-\epsilon_{h}q+F\left(q,X_{h}\right)+\frac{4\alpha}{r_{h}^{3}}+O(f),$ (28)
where $F$ depends on the four Horndeski functions $G_{i}$. As one can see, for an arbitrary value of $C$, the solution $\varphi^{\prime}$ exists with finite $X$ at horizons and our modified sGB behaves just like the Cubic Galileon. It must be stressed that our argument does not establish for which particular values of $C$ solutions regular in the whole intermediate spacetime actually exist. Determining when this construction is viable is beyond the scope of this letter.
One may still think that picking $C$ as in the static case (7) would give physics similar to the static case. This is not true. Notwithstanding whether such a solution would be regular, the background for $\phi$ would now be time dependent, $F$ would have to make a significant contribution to solve the two-horizon problem and, as a consequence, the perturbations would have very different dynamics. Observables such as radiated power and the quasi-normal-mode spectrum would be strongly affected.
VI Discussion & Conclusions
Let us take a different look at what has happened here: it is instructive to discuss shift-symmetric theories like sGB in terms of the dynamics of the corresponding shift charges – the equation of motion for the scalar (3) is a conservation equation for the shift-charge current, $J^{\mu}$ Pujolas et al. (2011), with $J^{0}$ the shift-charge density and $J^{i}$ is the spatial current giving the motion of these charges. The static sGB solution then corresponds to a static radial current (5) with flux $C$ across any surface enclosing the horizon. [^2] We stress that, in the shift-charge picture, $C$ is a flux of conserved shift charges the density of which is given by $J^{0}$, not a scalar charge as it is usually referred to. For this static configuration to be maintained, the charges must leave the spacetime at the rate they enter, and they are able to do so at spatial infinity. Keeping $X$ finite on a horizon means that the horizon must produce or absorb a particular flux $C$ of charges given by (7). Introducing a second horizon, e.g. cosmology, gives a different flux requirement. The two are in general incompatible, so the charges will begin to amass in the intermediate spacetime between the horizons breaking the staticity assumption of the solution.
Cosmology introduces time dependence to the scalar, with an attractor solution $q=q_{c}$ (20). This is also a vacuum solution with $J^{\mu}=0$ to which any black hole would have to match. This requires stationary time-dependent hair (16). However, the requirement of finiteness on both horizons requires a $q$ grossly different from $q_{c}$ and therefore a non-zero shift-charge density. This cannot be matched to a homogeneous cosmology.
We have not found any stationary solutions in sGB under our assumptions, but it may well be that the evolution away from it is slow and therefore one might be able to use the non-evolving $q=q_{c}$ solution with $C$ given by (16) for sufficiently long to study e.g. binary inspirals. Even so the static solution is not a good approximation since the cosmological background provides a medium which will affect the dynamics.
Alternatively, the introduction of (an) extra Horndeski terms resolves the inconsistency of the requirements put on the solution. The presence of $X$ in Eq. (28) means the theory is less rigid and allows for any choice of flux $C$ across any horizon even in the presence of the $q_{c}$ compatible with the cosmological vacuum needed to be reached at large distances. This is of course not sufficient that guarantee that such a solution exists – it only provides the boundary conditions on the two horizons which must then be connected through a non-singular source-free solution. This underlines how spectacularly the static sGB solution in cosmology fails – it cannot even support the required boundary conditions.
The price of this well-behaved alternative is that the hair which is necessary once a horizon appears in sGB Eq. (7) is no longer the same upon the introduction of any generic additional Horndeski operator. In any such good solution, the background that this hair gives for the fluctuations will be drastically different to the static sGB solution and we expect will strongly affect the predictions of any physical process. We leave the construction of viable configuration and the study of its properties to a separate work.
Acknowledgements.
VII Acknowledgments
The authors would like to thank Christos Charmousis, Gilles Esposito-Farèse and Nicolas Lecoeur for helpful discussions. E.B. acknowledges the support of ANR grant StronG (ANR-22-CE31-0015-01). The work of L.G.T. was supported by the European Union (Grant No. 101063210). I.S. is supported by the Czech Science Foundation (GAČR) project PreCOG (Grant No. 24-10780S). E.B. and L.G.T. would like to thank the hospitality of the Centro de Ciencias de Benasque Pedro Pascual during the initial stages of this work.